OpenStudy (anonymous):

A particle is moving with a constant speed; at time t=3 it has velocity <5,6> and acceleration . For what c is this possible?

OpenStudy (anonymous):

if a particle is moving with a constant speed then its acceleration = 0 but i dont know the notation you used <5,6> and <c,-4>

OpenStudy (anonymous):

HINT: That the speed is constant means that ||v(t)||=const or equivalently v(t)⋅v(t)=const. Differentiate both sides of this equality and use the properties of differentiation of vector functions (Stewart, page 566) and the properties of the dot product (Stewart, page 531) to obtain a relationship between the vectors v(t) and a(t)=v′(t). Use that relationship to find c.

OpenStudy (anonymous):

v(t) * v(t) = const d/dt [ v(t) * v(t) ] = d/dt [const] 2 * a(t) * v(t) = 0 it means that a(t) is perpendicular to v(t) so <5,6>* <c,−4> = 0 5c -24 = 0 c = 4.8